What will be the time required to half the amplitude of charge stored by a capacitor and time for the energy stored to be halved in a series LRC circuit with L=40H, R=0.03 ohms and C=4.8F?


1
Expert's answer
2021-11-16T13:47:27-0500

We know that

Charge

"q=q_0(1-e^{\frac{-t}{RC}})"

Gives

"q=\frac{q_0}{2}"

"\frac{q_0}{2}=q_0(1-e^{\frac{-t}{RC}})"

"\frac{1}{2}=(1-e^{\frac{-t}{RC}})"

"-\frac{1}{2}=e^\frac{-t}{RC}\\2=e^\frac{t}{RC}"

"t=RCln2"

Put value


"t=0.03\times4.8\times ln2=0.0998Sec"